OpenStudy (anonymous):

Could someone please teach me how to write a half life function from the function ln y = 0.18x + 0.3

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

if you want to solve for y, then you raise both sides as powers where the base is 'e' to get this \[\large \ln(y) = 0.18x + 0.3\] \[\large e^{\ln(y)} = e^{0.18x + 0.3}\] \[\large y = e^{0.18x + 0.3}\] Hopefully that's what you're looking for.